# Data-driven Quantum Dynamical Embedding Method for Long-term Prediction on Near-term Quantum Computers

### v3

@@ III.3 Experimental demonstration on a superconducting quantum processor
To further elaborate, we consider executing quantum circuits with grid inputs $m^{i},x^{i}$ in the state space $m,x\in[-1,1]$. The discrepancy between the ideal outputs, $(m^{*},x^{*})=\operatorname{QDE}(m,x)$, and actual outputs from the quantum processor without LECL, $(\hat{m},\hat{x})$, is quantified as $\epsilon_{m}=\hat{m}-m^{*}$ and $\epsilon_{x}=\hat{x}-x^{*}$. Subsequently, an LECL characterized by parameters $A$ and a bias vector $\mathbf{b}$ is applied to the experimental results within this state space. These parameters are optimized with Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm [59, 60, 61, 62], implemented in SCIPY [63], and fixed for the time series generation. Utilizing these optimized parameters, we construct a $25$-step time sequence that aligns with a single period of cosine-wave.
@@ III.5 Universality
In this section, we explore the universal applicability of the QDE in addressing a wide range of dynamical problems, with a specific focus on signal approximation.
III.5.1 Universality in the regime of signal approximation
We initiate our discussion by considering QDE’s ability to model a cosine-wave signal. The equations governing the QDE are shown in Eq. (16). Here, $\theta_{1}$ and $\theta_{2}$ are parameters that can be adjusted to model the signals $x(t)=A\cos(\omega t)=A\cos(\pi\Delta t)$ and $m(t)=A\sin(\omega t)=A\sin(\pi\Delta t)$, subject to amplitude $A<1/2$, $t\in\mathbb{N}$ and time step length $\Delta<1/20$. Let $\theta_{1}=-\theta<0$ and $\theta_{2}=\theta>0$, then we have the following lemmas.
@@ III.5.1 Universality in the regime of signal approximation
III.5.2 Ability for learning from non-linear dynamics
$\ddot{x}-\varepsilon\dot{x}(1-\delta x^{2})+\omega^{2}x=0,$ (19)
Figure 7: Simulation of Rayleigh Dynamics Using a $(1,2)$-QDE. This diagram illustrates the dynamics of the Rayleigh system, focusing on displacement $x_{t}$ and the velocity $v_{t}$, which constitute the data register in the QDE. An additional qubit is allocated for the memory register.
Introducing the velocity $v$, Eq. (19) can be recast as a system:
$\begin{cases}\dot{x}=v,\\ \dot{v}=\varepsilon v(1-\delta x^{2})-\omega^{2}x.\end{cases}$ (20)
To learn the dynamics of Rayleigh equation, the QDE model employs two qubits as the data register and one qubit as the memory register. This demonstration utilizes the ansatz TIEA as described in Sec. II.4. For the Rayleigh system, we set the parameters at $\varepsilon=\pi$, $\delta=3$, and $\omega=\pi$, initializing the system with $\bm{x}_{0}=(0,0.01)$. Fig. 7 presents the results, depicting two panels that represents $x(t)$ and $v(t)$ over time. The QDE model not only replicates the dynamics of the Rayleigh system but also exhibits consistent performance, closely following the ideal trajectory throughout the simulation.
@@ IV Discussion
Distinct from existing quantum-based methods for time series prediction, our method offers a practical solution for near-term quantum computers. It addresses the significant challenge of quantum decoherence, a common obstacle in deep quantum circuits, by employing the QDE within a fixed circuit depth. This design ensures that the circuit complexity does not increase with the length of the time series, making QDE particularly suited for current NISQ devices. Through numerical simulations and practical experiments, our method has proven effective in both long-term predicting capability and noise-resilient in practical scenarios.
Looking forward, the QDE lays a robust foundation for future explorations into the complex behaviors of systems with shallow quantum circuits. This paper paves the way to merge quantum time series algorithms with the study of dynamic systems, stepping into a new avenue for understanding the physical characteristics inherent to time series prediction.
@@ V Acknowledgements
This work was supported by National Key Research and Development Program of China (Grant No. 2023YFB4502500) and National Natural Science Foundation of China (Grant No. 12034018).
@@ I Quantum Reservoir Computing
where the disorder strength $D_{i}$ and the coupling strength $J_{ij}$ are randomly selected from uniform distributions in the interval $[-W,W]$ and the interval $[-J_{s},J_{s}]$, respectively. The transverse strength $h$ and the disorder bound $W$ will be expressed in units of coupling bound $J_{s}$ and $J_{s}=1$ for convenience. All coefficients remain fixed during training as shown in [21, 25]. Time multiplexing is employed to enhance learning, dividing the time interval $\tau$ into $V$ equal subintervals to increase the number of virtual nodes. The quantum state after $v$ subintervals’ evolution is given by:
@@
where $\mathbf{w}$, $\mathbf{y}$ and $\mathbf{X}$ are matrix forms corresponding to $w_{i}$, $\bm{y}_{t}$ and $x_{ki}$, respectively.The main difference between QRC and our QDE method lies in the placement of training parameters. The QDE method has parameters within the circuit, whereas the QRC model’s parameters are in the linear regression coefficients.
In our demonstrations, the state injection function is chosen as $\rho(s_{t})=\left|\Phi(s_{t})\right\rangle\left\langle\Phi(s_{t})\right|$, where $\left|\Phi(s_{t})\right\rangle=\sqrt{(1-s_{t})/2}\left|0\right\rangle+\sqrt{(1+s_{t})/2}\left|1\right\rangle$, considering the signal is one-dimensional and rescaled to the range $[-1,1]$.
@@ II Comparision of QDE with QRC
Figure S5: Comparison of QDE with QRC on Periodic Signal Prediction. (a) Error map for different values of the Hamiltonian hyperparameters of QRC, with the transverse field $h$ and the disorder bound $W$ varying logarithmically in the range $[10^{-2},10^{2}]$. Results are averaged over 100 realizations. (b) and (c) With $h/J_{s}$ ($W/J_{s}$) fixed at the minimum error from map (a), the other parameter $W/J_{s}$ ($h/J_{s}$) is varied. The average results are shown with the minimum to maximum error range (grey shadows) and standard deviation (red shadows).
@@
III Training Procedures
@@
IV Quantum Chip Information
The experiment presented in the main text was executed on the Origin “Wukong” superconducting quantum computer. This platform is composed of $72$ transmon qubits arranged in a two-dimensional array, with each qubit individually controlled via an XY line and Z control. The coupling strength of these qubits is adjustable from 00 to $80$ MHz, and the two-qubit CZ gate operates at a strength of $50$ MHz. The base frequencies of these transmon qubits are arranged in alternating high and low, the high frequency qubits have a sweetspot frequency of about $5$ GHz and the low are $4.5$ GHz.
@@ IV Quantum Chip Information
V Additional experiments for long-term time series prediction
Here, we demonstrate our results for single-mode cosine-wave signal and composite periodic signal for completeness, as shown in Fig. S8 and Fig. S9. The circuit architecture, the training and predicting length, and the noisy model are the same in the main text of Sec. II and Sec. III. The consistency with the theoretical curve is kept even after 1000 steps’ evolution. The decent performance of these tasks manifests the noise resistance of the proposed QDE model and the capability for capturing long-term features.
@@ V Additional experiments for long-term time series prediction
VI Detailed proofs
@@
Applying Lem. 1 and Lem. 2 with the initial conditions $m(0)=\hat{m}(0)$ and $x(0)=\hat{x}(0)$, and setting $-\theta_{1}=\theta_{2}=\pi\Delta$, we obtain:
@@
We can choose $c>b$ as the end of new interval $[a,c]$, which can be rescaled to $[a^{\prime},c^{\prime}]$ where $c^{\prime}-a^{\prime}=2\pi$. Then from Cor. 2 in Sec.1 of [65], there exists a trigonometric polynomial of the form
@@
for any $t^{\prime}$ in $[a^{\prime},c^{\prime}]$. The polynomial $\sigma_{J}(t^{\prime})$ can then be approximated using a $K$-channel QDE as described in Thm. 2, where in this case $K=2J+1$. ∎

### v4

@@ III.3 Experimental demonstration on a superconducting quantum processor
[…]
@@ III.5 Universality
In this section, we explore the applicability of the QDE framework in approximating composite cosine-wave signals, which serve as a foundational step toward universality. The extension of QDE to model more complex dynamical systems remains an open question and a subject for future research. We begin by examining QDE’s capability to model cosine-wave signals, governed by the equations in Eq. (16).
Here, $\theta_{1}$ and $\theta_{2}$ are parameters that can be adjusted to model the signals $x(t)=A\cos(\omega t)=A\cos(\pi\Delta t)$ and $m(t)=A\sin(\omega t)=A\sin(\pi\Delta t)$, subject to amplitude $A<1/2$, $t\in\mathbb{N}$ and time step length $\Delta<1/20$. Let $\theta_{1}=-\theta<0$ and $\theta_{2}=\theta>0$, then we have the following lemmas.
@@ IV Discussion
Distinct from existing quantum-based methods for time series prediction, our method offers a practical solution for near-term quantum computers. It addresses the significant challenge of quantum decoherence, a common obstacle in deep quantum circuits, by employing the QDE within a fixed circuit depth. Notably, the ansatz depth $C$ is significantly smaller than that required by full-sequence encoding methods, such as QRC, which scale as $C\times L$. Since the ansatz can be tailored to specific hardware constraints, QDE is particularly well-suited for current NISQ devices. This work represents an exploration of new effective strategies for temporal series processing, along with an acknowledgment of their limitations.
Through numerical simulations and practical experiments, our method demonstrates both long-term prediction capability and noise resilience in realistic scenarios. However, our simulations also reveal inherent optimization challenges, including the presence of local minima, the increased instability as the quantum system scales, and the barren plateau when the circuit depth of a single QDE block is sufficiently large. Additionally, the current methods for data extraction and injection within QDE do not preserve the full entanglement structure through each time step. This leads to inevitable information loss compared to approaches like the QRC model, which employs a quantum reservoir to maintain memory. The numerical results, while highlighting the limitations of this approach, raise important questions about whether both this method and previous ones offer quantum advantages in the domain of temporal processing, and how their classical simulability is affected given the reduced quantum resources required for the corresponding tasks. Addressing these questions is beyond the scope of the present work and will be the focus of our future researches.
Looking forward, the QDE framework provides a promising foundation for future investigations into leveraging quantum circuits with smaller depth to capture complex system behaviors. While our results demonstrate encouraging progress in integrating quantum time series algorithms with dynamical system analysis, further research is necessary to overcome current limitations and fully understand the physical characteristics inherent in long-term time series prediction.
@@ V Acknowledgements
We thank Song-Xin Qu for his valuable assistance with coding techniques and technical support. This work was supported by National Key Research and Development Program of China (Grant No. 2023YFB4502500) and National Natural Science Foundation of China (Grant No. 12034018).
@@ I Quantum Reservoir Computing
where the disorder strength $D_{i}$ and the coupling strength $J_{ij}$ are randomly selected from uniform distributions in the interval [W,W][-W,W] and the interval $J_{s},J_{s}]$, respectively. The transverse strength $h$ and the disorder bound $W$ will be expressed in units of coupling bound $J_{s}$ and $J_{s}=1$ for convenience. All coefficients remain fixed during training as shown in [22, 26]. Time multiplexing is employed to enhance learning, dividing the time interval $\tau$ into $V$ equal subintervals to increase the number of virtual nodes. The quantum state after $v$ subintervals’ evolution is given by:
@@
where $\mathbf{w}$, $\mathbf{y}$ and $\mathbf{X}$ are matrix forms corresponding to $w_{i}$, $\bm{y}_{t}$ and $x_{ki}$, respectively.The main difference between QRC and our QDE method lies in the placement of training parameters. The QDE method has parameters within the circuit, whereas the QRC model’s parameters are in the linear regression coefficients, with number of $(2n_{q}-1)\times V+1$.
In our demonstrations, the state injection function is chosen as $\rho(s_{t})=\left|\Phi(s_{t})\right\rangle\left\langle\Phi(s_{t})\right|$, where $\left|\Phi(s_{t})\right\rangle=\sqrt{(1-s_{t})/2}\left|0\right\rangle+\sqrt{(1+s_{t})/2}\left|1\right\rangle$, considering the signal is one-dimensional and rescaled to the range [1,1][-1,1].
@@ II Comparision of QDE with QRC
Figure S5: Comparison of QDE with QRC on Periodic Signal Prediction. (a) Error map for different values of the Hamiltonian hyperparameters of QRC, with the transverse field $h$ and the disorder bound $W$ varying logarithmically in the range $10^{-2},10^{2}]$. Results are averaged over 100 realizations. (b) and (c) With $h/J_{s}$ ($W/J_{s}$) fixed at the minimum error from map (a), the other parameter $W/J_{s}$ ($h/J_{s}$) is varied. The average results are shown with the minimum to maximum error range (grey shadows) and standard deviation (red shadows).
@@
III Ability for learning from non-linear dynamics
$\ddot{x}-\varepsilon\dot{x}(1-\delta x^{2})+\omega^{2}x=0,$ (S5)
Figure S7: Simulation of Rayleigh Dynamics Using a $(1,2)$-QDE. This diagram illustrates the dynamics of the Rayleigh system, focusing on displacement $x_{t}$ and the velocity $v_{t}$, which constitute the data register in the QDE. An additional qubit is allocated for the memory register.
Introducing the velocity $v$, Eq. (S5) can be recast as a system:
$\begin{cases}\dot{x}=v,\\ \dot{v}=\varepsilon v(1-\delta x^{2})-\omega^{2}x.\end{cases}$ (S6)
To learn the dynamics of Rayleigh equation, the QDE model employs two qubits as the data register and one qubit as the memory register. This demonstration utilizes the ansatz TIEA as described in Sec. II.4. For the Rayleigh system, we set the parameters at $\varepsilon=\pi$, $\delta=3$, and $\omega=\pi$, initializing the system with $\bm{x}_{0}=(0,0.01)$. Fig. S7 presents the results, depicting two panels that represents $x(t)$ and $v(t)$ over time. The QDE model not only replicates the dynamics of the Rayleigh system but also exhibits consistent performance, closely following the ideal trajectory throughout the simulation.
IV Training Procedures
@@
V Quantum Chip Information
The experiment presented in the main text was executed on the Origin “Wukong” superconducting quantum computer. This platform is composed of $72$ transmon qubits arranged in a two-dimensional array, with each qubit individually controlled via an XY line and Z control. The coupling strength of these qubits is adjustable from 0 to $80$ MHz, and the two-qubit CZ gate operates at a strength of $50$ MHz. The base frequencies of these transmon qubits are arranged in alternating high and low, the high frequency qubits have a sweetspot frequency of about $5$ GHz and the low are $4.5$ GHz.
@@ V Quantum Chip Information
VI Additional experiments for long-term time series prediction
Here, we demonstrate our results for single-mode cosine-wave signal and composite periodic signal for completeness, as shown in Fig. S9 and Fig. S10. The circuit architecture, the training and predicting length, and the noisy model are the same in the main text of Sec. II and Sec. III. The predictions remain consistent with the theoretical curves even after 1000 steps of evolution, demonstrating the QDE model’s noise resilience and its ability to capture long-term features effectively.
We note that the long-term test on the NARMA2 signal is conducted in a noise-free setting as shown in Fig. S11. Unlike the composite cosine-wave signals, which incorporate noise mitigation methods under depolarizing and phase damping noise, the NARMA2 results serve as a comparison rather than a demonstration of noise robustness. This choice does not imply that the method is unsuitable for noisy NARMA2 tasks.
@@ VI Additional experiments for long-term time series prediction
Figure S11: Long-term Prediction on Noise-Free NARMA2 Signal.
VII Detailed proofs
@@
Applying Lem. 1 and Lem. 2 with the initial conditions $m(0)=\hat{m}(0)$ and $x(0)=\hat{x}(0)$, and setting $\theta_{1}=\theta_{2}=\pi\Delta$, we obtain:
@@
We can choose $c>b$ as the end of new interval $[a,c]$, which can be rescaled to $a^{\prime},c^{\prime}]$ where $c^{\prime}-a^{\prime}=2\pi$. Then from Cor. 2 in Sec.1 of [72], there exists a trigonometric polynomial of the form
@@
for any $t^{\prime}$ in $a^{\prime},c^{\prime}]$. The polynomial $\sigma_{J}(t^{\prime})$ can then be approximated using a $K$-channel QDE as described in Thm. 2, where in this case $K=2J+1$. ∎